QUES : - find the volume of the water dis placed and position of centre of buoyancy for a wooden block of width 2.5m and of Depth 1.5m when it floats horizontally in water the density of the all wooden block is 650kg/m3 and its length 6m.
1
Expert's answer
2020-04-27T10:14:45-0400
"m_w = d_wV"

"V_{block} = V_{water}"

"V_{block} = 2.5m*1.5m*6m = 22.5m^3"

"P_{inblock} = m_{block}g - m_{water}g"

"P_{inblock} = 22.5*650*10 - 22500*10 = -78750N"

"P_{centre} = 146250N-78750N = 67500N"